How do you identify the slope and y-intercept of y-16=4x.
Answer by macston(5194)

In slope intercept form:
y=mx+b, m=slope and b=y value of y-intercept
Put the given equation in slope intercept form:
y-16=4x Add 16 to each side.
y=4x+16 in this equation, m=4 and b=16.
The slope=m=4, y-intercept=b=16